# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

### v1

@@ 4.5 Role of Fact-Level Detection in Hallucination Correction
One potential application of hallucination detection methods is their use in correcting hallucinated responses. In this experiment, we investigate the effectiveness of our fact-level detection approach in enhancing hallucination correction and compare the results with those obtained using sentence-level detection and a baseline method. Each of the three tested approaches uses different input for the LLM: (1) Baseline: the original prompt and the generated response (Prompt 7). (2) Sentence-level: the original prompt, the generated response, and a list of hallucinated sentences (Prompt 8). (3) Fact-level: the original prompt, the generated response, and a list of hallucinated facts (Prompt 9).
The original prompt is the one used during the creation of the dataset: "This is a Wikipedia passage about {concept_name}:". We instructed the LLM to return a list of sentences, allowing it to correct each sentence or leave it unchanged if no hallucinations were detected. We obtained the lists of incorrect sentences/facts using the best variants of models (see Section 5.1) with thresholds that achieved the highest F1-scores on the dataset ($0.3$ for FactSelfCheck and $0.75$ for SelfCheckGPT).
Subsequently, we evaluated the factuality of the corrected responses using the LLM-as-judge approach Zheng et al. (2023) using Prompt 10. For each corrected sentence, we provided the external biography from Wikipedia. We instructed LLM-as-judge to return ’yes’ if the source supported the sentence, ’no’ if it was not, or ’refused’ if the LLM declined to correct the sentence (e.g., due to insufficient knowledge). We then categorized the responses into three labels: ’factual’, ’non-factual’, ’refused’.
As mentioned in Section 4.2, we utilized GPT-4o for this experiment instead of Llama-3.1-70B-Instruct (used in detection). This choice was motivated by the challenging nature of the correction task – the model needs to correct hallucinations using only its internal knowledge, without access to external references. While the model knows hallucinated parts, it must rely on its knowledge to determine the correct information.
While the described correction method is not our main contribution, we used it to study the potential benefits of fact-level detection. Although the correction method employed here may not be the most sophisticated, the key takeaway is the observed difference in performance.
@@ 5 Results
This section presents the results of the experiments described in the previous section. The results of the fact-level evaluation are given in Appendix B.
5.1 Evaluation of Sentence-Level and Passage-Level Hallucination Detection
@@
SCGPT (Unigram-max) - 85.63
SCGPT (QA) - 84.26
SCGPT (BERTScore) - 81.96
@@
FSC-KG (Freq.-based) max 88.48
FSC-KG (Freq.-based) mean 88.25
@@
Table 1: Results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores.
Method Agg. Pear. Spear.
@@
SCGPT (Prompt) - 83.12 78.65
SCGPT (NLI) - 74.14 73.78
SCGPT (Unigram-max) - 64.71 64.91
SCGPT (QA) - 61.07 59.29
SCGPT (BERTScore) - 58.18 55.90
@@
FSC-Text max 79.69 76.68
FSC-KG (LLM-based) max 77.30 75.37
FSC-Text mean 77.03 75.89
FSC-KG (LLM-based) mean 73.92 72.28
FSC-KG (Freq.-based) mean 71.50 70.45
FSC-KG (Freq.-based) max 64.05 66.09
@@
Table 2: Results on the passage-level hallucination detection task. Comparison of sentence-level and fact-level methods based on Pearson and Spearman correlation scores. Results are sorted by the Pearson correlation score. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores. As mentioned in Section 4, passage-level scores were calculated using the mean of sentence-level scores.
@@
Tables 1 and 2 present a comparative analysis of our method against SelfCheckGPT (SCGPT). In the sentence-level evaluation, FactSelfCheck-Text utilizing $max$ as an aggregation function achieves an AUC-PR score of $92.45$. It demonstrates that our approach is comparable in performance to the leading SCGPT variants – Prompt ($93.60$) and NLI ($92.50$). Notably, while our method operates at a more granular level, it maintains competitive performance with a marginal decrease of $1.2\%$ compared to the best SCGPT. Figure 3 illustrates the precision-recall trade-off across various detection thresholds.
In the passage-level evaluation, our method achieves Pearson and Spearman correlation coefficients of $79.69$ and $76.68$, respectively. Although these results are lower than SCGPT (Prompt) ($83.12$ and $78.65$), they demonstrate the effective scalability of the fact-level detection to the sentence-level and passage-level detections.
@@
Among the FactSelfCheck (FSC) variants, FSC-Text, which performs the direct comparison between facts and samples, consistently outperforms FSC-KG, which relies on knowledge graph comparisons. FSC-Text offers computational advantages by eliminating the knowledge graph extraction step. However, FSC-KG may be more computationally efficient for processing longer samples with a lower fact density because of reduced token usage. Both variants perform better than the Frequency-based FSC, which employs a simple fact occurrence counting mechanism. This performance differential highlights the benefits of incorporating semantic matching and knowledge graph reasoning in hallucination detection. Nevertheless, the frequency-based method’s computational efficiency may make it suitable for applications where its reduced accuracy is acceptable. Regarding aggregation functions, $max$ consistently yields better results than $mean$, as it is more sensitive to the presence of hallucinated facts in the text.
@@
An interesting side observation is that our reproduced SCGPT on Llama-3.1-70B-Instruct marginally surpassed the original implementation using GPT-3.5-turbo ($93.60$ vs $93.42$  [^5] The scores on GPT-3.5-turbo were obtained from the original paper Manakul et al. (2023).). Furthermore, for passage-level detection, Llama achieved a better Pearson correlation coefficient than the GPT-3.5-turbo ($83.12$ vs. $78.32$ 5), while the Spearman correlation coefficient remained at a similar level ($78.65$ vs. $78.30$ 5). This indicates that contemporary open-source models can exceed the performance of previous closed models.
5.2 Effect of Sample Size on Detection Performance
Figure 2: Impact of the number of samples on sentence-level detection task.
Figures 2 and 4 illustrate the relationship between the number of samples and detection performance for both FactSelfCheck variants and SelfCheckGPT (Prompt). The results show that FactSelfCheck exhibits similar behavior to SelfCheckGPT regarding sample requirements. The performance improves dramatically with up to $5$ samples, after which the improvement curve flattens. While additional samples continue to yield benefits, these improvements become incremental, with modest gains observed up to $20$ samples.
@@ 5.3 Role of Fact-Level Detection in Hallucination Correction
Table 3 presents the results of our hallucination correction experiment. The fact-level approach shows substantial improvements over the baseline and sentence-level methods. We observed a $35\%$ increase in factual content and $12\%$ reduction in non-factual content compared to the baseline. In contrast, the sentence-level detection achieves only modest improvements of $8\%$ and $4\%$, respectively, indicating that pointing out hallucinations at the fact level enables more effective corrections and underscoring the importance of our study and contributions.
@@ 6 Conclusion
Our study introduces FactSelfCheck, a novel approach to hallucination detection that operates at the fact level rather than the sentence level or passage level. The evaluation results demonstrate that this granular approach achieves competitive performance with methods like SelfCheckGPT, with the AUC-PR score of $92.45$ on sentence-level detection compared to SelfCheckGPT (Prompt)’s $93.60$, while offering more detailed insights into the nature of hallucinations.
Our fact-level approach improves hallucination correction. Providing the correction system with a list of incorrect facts rather than incorrect sentences allows for more effective corrections. Compared to baseline, we achieved $35\%$ increase in factual content and $12\%$ reduction in non-factual content. In contrast, the sentence-level approach achieved a $8\%$ increase and $4\%$ reduction, respectively. It underscores the importance of fine-grained hallucination detection in enhancing the reliability and trustworthiness of systems that use LLMs, particularly in applications requiring high factual accuracy.
@@
Our study faces two primary limitations. First, we are constrained by the availability of suitable datasets. Currently, the WikiBio GPT-3 Hallucination Dataset, to the best of our knowledge, is the only dataset designed for evaluating sampling-based methods, and it only provides sentence-level and passage-level annotations. It forced us to evaluate our method through aggregation rather than directly assessing our fact-level detection capabilities. Second, while the granular approach of FactSelfCheck justifies its increased complexity, the multiple LLM-based steps make it more computationally intensive compared to more straightforward methods like SelfCheckGPT.
Several promising directions could address these limitations. An important step would be creating new datasets with fact-level hallucination annotations, enabling direct evaluation of our method’s core capabilities. Additionally, we see significant potential for improving computational efficiency. Our current prompt engineering was largely empirical and not optimized for token usage. Future work could focus on reducing prompt lengths and merging steps, such as simultaneously assessing support for multiple facts or merging the KG extraction steps. Finally, incorporating external knowledge graphs could enhance fact verification while potentially reducing our reliance on stochastic sampling.
@@
Like all machine learning methods, FactSelfCheck can produce false positives and false negatives. Therefore, it should not completely replace human verification of factual correctness in LLM responses. Additionally, since the fact-level annotations (Appendix B) are generated by an LLM, there remains a risk of fact mislabeling.
@@ Acknowledgments
Appendix A Additional Figures for Results
Figure 3 depicts the precision-recall curve for sentence-level detection (see Section 5.1). Figure 4 shows the impact of the number of samples on the passage-level detection task (see Section 5.2).
Figure 3: Precision-recall curve for the sentence-level detection task.
Figure 4: Impact of the number of samples on passage-level detection task.
Appendix B Fact-Level Detection Evaluation
While the evaluation in Section 5.1 focused on the sentence-level and passage-level detection due to dataset constraints, we present a complementary evaluation at the fact level here. Since this evaluation is not based on human annotations, we have placed it in the appendix.
@@ Appendix B Fact-Level Detection Evaluation
To enable fact-level evaluation, we annotated facts from the extracted response knowledge graph ($KG_{p}$) using the LLM-as-judge approach Zheng et al. (2023). This approach annotated whether each fact is supported by the external Wikipedia biography (Prompt 11). As a result, we obtained $5488$ binary annotated facts.
@@
We evaluated all variants of FactSelfCheck using the fact-level score $\mathcal{H}_{\text{fact}}$. For comparison with SelfCheckGPT, which provides only sentence-level granularity, we derived fact-level scores by averaging the sentence-level scores across all sentences containing each fact.
@@ Results
Appendix C Passage-level Score: Size of Knowledge Graphs Extracted from Samples
As a side study, we examined the effectiveness of scoring hallucinations at the passage level based on the size of knowledge graphs extracted from samples.
Score Definition
We propose using the average size of restricted knowledge graphs extracted from samples as a passage-level score. Restriction involves keeping only facts where all entities and relations are present in $\mathcal{E}_{S}$ or $\mathcal{R}_{S}$, respectively. This results in $KG_{s}^{\prime}$ defined as:

### v2

@@ 4.5 Role of Fact-Level Detection in Hallucination Correction
One potential application of hallucination detection methods is their use in correcting hallucinated responses. In this experiment, we investigate the effectiveness of our fact-level detection approach in enhancing hallucination correction and compare the results with those obtained using sentence-level detection and a baseline method. Each of the three tested approaches uses different input for the LLM: (1) Baseline: the original prompt and the generated response. (2) Sentence-level: the original prompt, the generated response, and a list of hallucinated sentences. (3) Fact-level: the original prompt, the generated response, and a list of hallucinated facts.
As only the WikiBio dataset provides reference in form of the real Wikipedia biography, and the FavaMultiSamples dataset does not, we conducted this experiment only on WikiBio. The original prompt is the one used during the creation of the dataset: "This is a Wikipedia passage about {concept_name}:". We instructed the LLM to return a list of sentences, allowing it to correct each sentence or leave it unchanged if no hallucinations were detected. We obtained the lists of incorrect sentences/facts using the best variants of models on this dataset (see Section 5.1) with thresholds that achieved the highest F1-scores on the dataset ($0.3$ for FactSelfCheck-Text and $0.75$ for SelfCheckGPT (Prompt)).
Subsequently, we evaluated the factuality of the corrected responses using the LLM-as-judge approach Zheng et al. (2023). For each corrected sentence, we provided the external biography from Wikipedia as a reference. The reliability of this approach is high, and commonly used in the literature Thakur et al. (2025). We instructed LLM-as-judge to return ’yes’ if the source supported the sentence, ’no’ if it was not, or ’refused’ if the LLM declined to correct the sentence (e.g., due to insufficient knowledge). We then categorized the responses into three labels: ’factual’, ’non-factual’, ’refused’. As mentioned in Section 4.3, we utilized GPT-4o for correction and judging instead of Llama-3.1-70B-Instruct (used in detection). This choice was motivated by the challenging nature of the correction task – the model needs to correct hallucinations using only its internal knowledge, without access to external references. While the model knows hallucinated parts, it must rely on its knowledge to determine the correct information.
While the method of correction described here is not our main contribution, we used it to study the potential benefits of fact-level detection. Although the correction method employed here may not be the most sophisticated, the key takeaway is the observed difference in performance.
@@ 5 Results
This section presents the results of the experiments described in Section 4. The results of the fact-level evaluation are presented in Appendix C.
5.1 Sentence-Level Detection
@@
AttentionScore (Relative) - 83.85
$\mathcal{H}$) - 82.56
[…]
$\mathcal{H}$) - 79.02
[…]
AttentionScore (Absolute) - 77.95
@@
FSC-KG (Frequency-based) max 88.48
FSC-KG (Frequency-based) mean 88.25
@@
Table 1: WikiBio: Results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores.
Method Agg. AUC-PR
@@
SCGPT (Prompt) - 46.91
SCGPT (NLI) - 32.58
$\mathcal{H}$) - 28.22
[…]
AttentionScore (Relative) - 24.17
$\mathcal{H}$) - 23.80
[…]
AttentionScore (Absolute) - 22.19
RandomSentence - 21.70
@@
FSC-KG (Frequency-based) max 48.52
FSC-Text max 42.80
FSC-KG (LLM-based) max 40.63
FSC-Text mean 37.13
FSC-KG (Frequency-based) mean 36.16
FSC-KG (LLM-based) mean 35.81
RandomFact mean 21.22
@@
Table 2: FavaMultiSamples: Results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores.
@@
Tables 1 and 2 present a comparative analysis of our method against baselines. For WikiBio, FactSelfCheck-Text utilizing $max$ as an aggregation function achieves an AUC-PR score of $92.45$. It demonstrates that our approach is comparable in performance to the leading SelfCheckGPT (SCGPT) variants – Prompt ($93.60$) and NLI ($92.50$). Notably, while our method operates at a more granular level, it maintains competitive performance with a marginal decrease of $1.2\%$ compared to the best SCGPT. It is important to note that comparing our method to SelfCheckGPT at the sentence level inherently disadvantages our approach, as we operate at a more granular level of analysis. Nevertheless, our method still achieves competitive performance despite this inherent challenge.
@@
For FavaMultiSamples, FactSelfCheck-KG (Frequency-based) with max aggregation achieves the highest AUC-PR score of $48.52$, outperforming all sentence-level baselines, including SelfCheckGPT (Prompt) at $46.91$. This result, together with the findings from WikiBio, highlights that the best-performing FactSelfCheck (FSC) variant depends on the dataset. On WikiBio, FSC-Text, which performs direct comparison between facts and samples, consistently achieves the highest hallucination AUC-PR ($92.45$), outperforming FSC-KG, which relies on knowledge graph comparisons. FSC-Text offers computational advantages by eliminating the knowledge graph extraction step. However, on FavaMultiSamples, the frequency-based FSC-KG variant surpasses both FSC-Text and LLM-based FSC-KG. These differences suggest that the optimal variant is influenced by dataset characteristics, such as fact density, sentence length, and text style (e.g., the prevalence of lists in FavaMultiSamples). FSC-KG may be more computationally efficient for longer samples with lower fact density due to reduced token usage, while FSC-Text is preferable for shorter or denser samples. Regarding aggregation functions, $max$ consistently yields better results than $mean$, as it is more sensitive to the presence of hallucinated facts in the text.
@@
An interesting side observation is that our reproduced SCGPT (Prompt) with Llama-3.1-70B-Instruct marginally surpassed the original implementation using GPT-3.5-turbo ($93.60$ vs $93.42$  [^7] The scores on GPT-3.5-turbo were obtained from the original paper Manakul et al. (2023).).
@@ 5.2 Role of Fact-Level Detection in Hallucination Correction
Table 3 presents the results of our hallucination correction experiment. The fact-level approach shows substantial improvements over the baseline and sentence-level methods. We observed a $35.5\%$ increase in factual content and $12.5\%$ reduction in non-factual content compared to the baseline. In contrast, the sentence-level detection achieves only modest improvements of $10.6\%$ and $4.8\%$, respectively, indicating that pointing out hallucinations at the fact level enables more effective corrections and underscoring the importance of our study and contributions.
@@ 6 Conclusion
We introduced FactSelfCheck, a novel approach to hallucination detection that operates at the fact-level rather than the sentence or passage level. Our evaluation demonstrates that this fine-grained approach achieves competitive performance with SelfCheckGPT, a less granular method. On the sentence-level detection task, FactSelfCheck achieves AUC-PR scores of $92.45$ and $48.52$ on WikiBio and FavaMultiSamples, respectively, compared to SelfCheckGPT’s $93.60$ and $46.91$. While maintaining comparable performance, our method provides more detailed insights into the nature of hallucinations.
Our fact-level approach improves hallucination correction. Providing the correction system with a list of incorrect facts rather than incorrect sentences allows for more effective corrections. Compared to baseline, we achieved $35.5\%$ increase in factual content and $12.5\%$ reduction in non-factual content. In contrast, the sentence-level approach achieved a $10.6\%$ increase and $4.8\%$ reduction, respectively. It underscores the importance of fine-grained hallucination detection in enhancing the reliability and trustworthiness of systems that use LLMs, particularly in applications requiring high factual accuracy. Finally, we contributed FavaMultiSamples, a novel dataset for evaluating sampling-based methods.
@@
Our study faces two primary limitations. First, we are constrained by the availability of suitable datasets. Although we contributed FavaMultiSamples, the second dataset for evaluating sampling-based methods, it is still a dataset with annotations at the sentence-level. The lack of datasets with long generated passages and annotations at the fact-level forced us to evaluate our method through aggregation rather than directly assessing our fact-level detection capabilities. Second, while the granular approach of FactSelfCheck justifies its increased complexity, the multiple LLM-based steps make it more computationally intensive compared to more straightforward methods like SelfCheckGPT.
Several promising directions could address these limitations. An important step would be creating new datasets with fact-level hallucination annotations, enabling direct evaluation of our method’s core capabilities. Additionally, we see significant potential for improving computational efficiency. Our current prompt engineering was largely empirical and not optimized for token usage. Future work could focus on reducing prompt lengths and merging steps, such as merging the KG extraction steps or simultaneously assessing support for multiple facts.
@@
Like all machine learning methods, FactSelfCheck can produce false positives and false negatives. Therefore, it should not completely replace human verification of factual correctness in LLM responses.
@@ Acknowledgments
Appendix A Fava-Sampling Dataset
We built FavaMultiSamples upon the FAVA dataset [^8] https://huggingface.co/datasets/fava-uw/fava-data license: CC-BY-4.0 developed by Mishra et al., 2024. The original FAVA dataset contains $460$ passages generated by GPT (gpt-3.5-turbo-0301) and Llama2-Chat-70B in response to diverse information-seeking prompts. Each passage was annotated by trained annotators for factual accuracy. For more details about the dataset construction and annotation process, please refer to the original paper.
To create FavaMultiSamples, we generated $20$ samples for each passage with the temperature of $1.0$, matching the sample settings used in the WikiBio dataset. We used the same models that produced the original responses, with one exception: since gpt-3.5-turbo-0301 is no longer available, we used gpt-3.5-turbo-1106, the most similar model available at the time. The FAVA dataset uses an HTML-like format for annotations, so we split each generated response into sentences and annotated them in binary format, where $1$ indicates a sentence containing a hallucination.
Appendix B Dataset Statistics
The statistics of the used datasets are presented in Table 4.
WikiBio FavaMulti-
Samples
# Passages 238 460
# Sentences 1908 5660
# Hall. sentences 1392 1228
# Fact. sentences 516 4432
% Hall. sentences 72.96 21.70
Avg. sent./passage 8.02 12.30
Avg. tok./passage 184.77 340.72
Avg. tok./sentence 23.48 30.30
Table 4: Statistics of the used datasets: WikiBio and FavaMultiSamples. We summarize the number of passages, sentences, tokens and annotation statistics. The number of tokens were calculated using appropriate tokenizers for each model.
Appendix C Fact-Level Detection Evaluation
While the evaluation in Section 5.1 focused on the sentence-level detection due to dataset constraints, we present a complementary evaluation at the fact level here. Since this evaluation is not based on human annotations, we have placed it in Appendix.
@@ Appendix C Fact-Level Detection Evaluation
To enable fact-level evaluation, we annotated facts from the extracted response knowledge graph ($KG_{p}$) using the LLM-as-judge approach Zheng et al. (2023). Since we pass the external Wikipedia biography to the LLM, the reliability of this approach is high, and commonly used in the literature Thakur et al. (2025). This approach annotated whether each fact is supported by the external Wikipedia biography. As a result, we obtained $5488$ binary annotated facts.
@@
We evaluated all variants of FactSelfCheck using the fact-level score $\mathcal{H}_{\text{fact}}$. For comparison with SelfCheckGPT we used the best performing variant – Prompt. As SelfCheckGPT provides only sentence-level granularity, we derived fact-level scores by averaging the sentence-level scores across all sentences containing each fact.
@@ Results
Comparing FactSelfCheck to SelfCheckGPT could be seen as unfair because SCGPT operates at a lower granularity than required by the evaluation task. However, this situation is analogous to the sentence-level evaluation presented in Section 5.1. In both cases, direct comparisons are not fully appropriate. The key difference is that, at the fact level, FactSelfCheck significantly outperforms SelfCheckGPT, while at the sentence level, FactSelfCheck remains competitive despite the unfavorable comparison.
Appendix D Additional Results
D.1 Sentence-Level Detection
Tables 6 and 7 provide a more comprehensive view of the FactSelfCheck variants against baselines on both datasets, including AUC-PR scores for detecting factual sentences (Factuality AUC-PR) in addition to hallucinated ones, and an average of these two. While Section 5.1 focused on hallucination detection, these extended results offer additional insights into our approach. On WikiBio, FSC-Text (max aggregation) shows comparable performance in hallucination detection with $92.45$ AUC-PR vs. $93.60$ for SCGPT (Prompt). While SCGPT (Prompt) achieves a higher factuality detection score ($74.30$ vs. $65.55$), our method performs well in identifying potential misinformation and contributes to effective factual verification.
On the FavaMultiSamples dataset, FSC-KG (Frequency-based, max aggregation) demonstrates high overall performance, achieving a factuality detection AUC-PR of $80.21$ alongside solid hallucination detection ($48.52$ AUC-PR compared to SCGPT Prompt’s $46.91$). This performance across metrics shows the adaptability of our approach to different datasets. While SCGPT (Prompt) achieves a higher average AUC-PR ($67.65$ vs. $64.36$) due to stronger factuality detection ($88.39$), our method provides balanced performance across both metrics, offering advantages for applications where hallucination detection is important.
Figures 2 and 3 illustrate the precision-recall curves for the sentence-level detection task on the WikiBio and FavaMultiSamples datasets, respectively. These curves show the performance across various thresholds for both hallucination and factuality detection.
Method Agg. AUC-PR
Hallucination Factuality Avg.
Sentence-level methods
SCGPT (Prompt) - 93.60 74.30 83.95
SCGPT (NLI) - 92.50 66.08 79.29
AttentionScore (Relative) - 83.85 51.62 67.74
$\mathcal{H}$) - 82.56 41.80 62.18